# Learning large softmax mixtures with warm start EM

### v1

@@ Abstract
Mixed multinomial logits are discrete mixtures introduced several decades ago to model the probability of choosing an attribute $x_{j}\in\mathbb{R}^{L}$ from $p$ possible candidates, in heterogeneous populations. The model has recently attracted attention in the AI literature, under the name softmax mixtures, where it is routinely used in the final layer of a neural network to map a large number $p$ of vectors in $\mathbb{R}^{L}$ to a probability vector. Despite its wide applicability and empirical success, statistically optimal estimators of the mixture parameters, obtained via algorithms whose running time scales polynomially in $L$, are not known. This paper provides a solution to this problem for contemporary applications, such as LLMs (Large Language Models), in which the mixture has a large number $p$ of support points, and the size $N$ of the sample observed from the mixture is also large. Our proposed estimator combines two classical estimators, obtained respectively via a method of moments (MoM) and the expectation-minimization (EM) algorithm. Although both estimator types have been studied, from a theoretical perspective, for Gaussian mixtures, no similar results exist for softmax mixtures for either procedure. We develop a new MoM parameter estimator based on latent moment estimation that is tailored to our model, and provide the first theoretical analysis for a MoM-based procedure in softmax mixtures. Although consistent, as $N,p\rightarrow\infty$, MoM for softmax mixtures can exhibit poor numerical performance, an empirical observation that is in line with those made for other mixture models. Nevertheless, as MoM is provably in a neighborhood of the target, it can be used as warm start for any iterative algorithm. We study in detail the EM algorithm, and provide its first theoretical analysis for softmax mixtures, extending the only other class of similar results, valid for Gaussian mixtures. Our final proposal for parameter estimation is the EM algorithm with a MoM warm start. In addition to leading to the desired parametric estimation rates, this combined procedure provides computational savings relative to the standard practice of selecting one of the outputs of multiple EM runs, each initialized at random. These facts are supported by our simulation studies. Concrete examples that substantiate the large applicability of the model are given in the Introduction.
@@ 1 Introduction
“Softmax mixtures” define a parametric mixture model in $\Delta_{p}$, the probability simplex in $\mathbb{R}^{p}$: for a known set of vectors $x_{1},\dots,x_{p}\in\mathbb{R}^{L}$, each component of a softmax mixture is obtained from $x_{1},\dots,x_{p}$ and an unknown parameter ${\bm{\theta}}\in\mathbb{R}^{L}$ via the softmax function:
$\mathsf{softmax}:\mathbb{R}^{p}\to\Delta_{p},\quad\quad\bigl{[}\mathsf{softmax}(x_{1}^{\top}{\bm{\theta}},\ldots,x_{p}^{\top}{\bm{\theta}})\bigr{]}_{j}=\frac{{\exp({x_{j}^{\top}{{\bm{\theta}}}})}}{\sum_{i=1}^{p}\exp({x_{i}^{\top}{{\bm{\theta}}}})},\quad\mbox{for each}\ j\in[p].$
For any integer $k\geq 1$, we write $[k]:=\{1,\ldots,k\}$. Widely used in statistical analyses of categorical data, where they are more commonly known as mixtures of multinomial logits, softmax mixtures have also seen a revival in modern AI applications, where the softmax function is routinely used as the final layer in a deep neural network, such as a large language model, to map a large collection of $p$ vectors in $\mathbb{R}^{L}$ to a probability distribution in $\Delta_{p}$.
The aim of this work is to provide a framework within which the parameters of such mixtures can be estimated and analyzed, with a focus on the regime where $p$ is large and $x_{1},\dots,x_{p}$ are random. To this end, it is convenient to view this model not as a mixture model on $\Delta_{p}$, but as a $p$-dimensional discrete conditional mixture model on $\mathbb{R}^{L}$, in a sense described precisely below.
One begins by observing random vectors in $\mathbb{R}^{L}$, for some $L<p$,
$X_{1},\dots,X_{p}\ {\mbox{i.i.d.}}\ \mu,$ (1)
where $\mu$ is a distribution on $\mathbb{R}^{L}$.
The softmax mixture model is a discrete mixture distribution which, conditionally on having observed $X_{1}=x_{1},\ldots,X_{p}=x_{p}$, is also supported on $x_{1},\ldots,x_{p}$. Each mixture component is a discrete distribution on these $p$ points, parametrized by some ${{\bm{\theta}}}\in\mathbb{R}^{L}$ via the softmax function
$\quad A_{{\bm{\theta}}}(x_{j}\mid x_{1},\ldots,x_{p}):=\bigl{[}\mathsf{softmax}(x_{1}^{\top}{\bm{\theta}},\ldots,x_{p}^{\top}{\bm{\theta}})\bigr{]}_{j}=\frac{{\exp({x_{j}^{\top}{{\bm{\theta}}}})}}{\sum_{i=1}^{p}\exp({x_{i}^{\top}{{\bm{\theta}}}})},\quad\mbox{for each}\ j\in[p].$ (2)
Then, the discrete mixture supported on $\{x_{1},\ldots,x_{p}\}$ with masses given by
$\pi_{{\bm{\omega}}}(x_{j}\mid x_{1},\ldots,x_{p}):=\sum_{k=1}^{K}\alpha_{k}A_{{\bm{\theta}}_{k}}(x_{j}\mid x_{1},\ldots,x_{p}),\qquad\mbox{for each}\ ~{}j\in[p],$ (3)
for distinct vectors ${\bm{\theta}}_{k}\in\mathbb{R}^{L}$, $k\in[K]$ and mixture vector ${\bm{\alpha}}:=(\alpha_{1},\ldots,\alpha_{K})^{\top}\in\Delta_{K}$, for some fixed integer $K$, is called a softmax mixture. We collect all model parameters in ${\bm{\omega}}:=({{\bm{\theta}}}_{1},\ldots,{{\bm{\theta}}}_{K},\bm{\alpha}),$ and denote the true value by ${\bm{\omega}}^{*}$.
Conditionally on having observed $x_{1},\ldots,x_{p}$, one then observes a sample from this mixture
$Y_{1},\ldots,Y_{N}\mid X_{1}=x_{1},\ldots,X_{p}=x_{p}\quad\text{i.i.d.}\ \pi_{{\bm{\omega}}^{*}}(\cdot\mid x_{1},\ldots,x_{p}),$ (4)
The goal of this paper is to provide and analyze estimators of ${\bm{\omega}}^{*}$ from samples given by (1) and (4), when $\mu=\mathcal{N}_{L}(0,\Sigma)$.
We note that we can interpret $Y_{1},\ldots,Y_{N}$ as being obtained by re-sampling from the collection $\{X_{1},\ldots,X_{p}\}$. While in standard re-sampling procedures, such as the bootstrap, the re-sampling probabilities are given, the new paradigm here is that the re-sampling probabilities $\pi_{{\bm{\omega}}^{*}}(x_{j}\mid x_{1},\ldots,x_{p})$ are unknown, and ${\bm{\omega}}^{*}$ needs to be estimated from the data.
@@
When $K=1$, the softmax mixture model reduces to what is known in the classical statistical literature as the conditional logit model (McFadden, 1974). Its usage and properties, when the features $x_{1},\ldots,x_{p}$ are not random vectors and $p$ is fixed, have been thoroughly studied, see McFadden (1974) and the literature review in Agresti (1990), Chapter 9, for instance. Much less is known about the case $K>1$, which has received very little attention in the mathematical statistics literature. Nevertheless, this model is widely used in applied contexts. We highlight a few prominent examples of softmax mixtures:
  • Discrete choice models. Softmax mixtures were introduced in the econometrics literature by Boyd and Mellman (1980) and Cardell and Dunbar (1980) under the name “mixed multinomial logits” to model the preference of a heterogeneous set of consumers for a set of mutually exclusive goods. In this application, each vector $X_{j}\in\mathbb{R}^{L}$ reflects the set of attributes of each of the $p$ different goods, while the vector ${{\bm{\theta}}}$ reflects a customer’s preferences for each attribute. The model posits that customers act via random utility maximization: the customer chooses good $j^{\star}\in[p]$ given by
    $j^{\star}=\operatorname*{argmax}_{j\in[p]}X_{j}^{\top}{{\bm{\theta}}}+\epsilon_{j}\,,$
    where $\epsilon_{1},\ldots,\epsilon_{p}$ are independent stochastic terms that reflect idiosyncratic variations in the consumer’s taste. When $\epsilon_{1},\ldots,\epsilon_{p}$ are chosen to have a Gumbel distribution, then (see, e.g. Yellott Jr, 1977)
    $\mathbb{P}\left\{j^{\star}=j\mid X_{1}=x_{1},\dots,X_{p}=x_{p}\right\}=\frac{\exp(x_{j}^{\top}{\bm{\theta}})}{\sum_{i=1}^{p}\exp(x_{i}^{\top}{\bm{\theta}})}\quad\quad\text{for each $j\in[p]$,}$
    so that a customer with preference vector $\bm{{\bm{\theta}}}$ chooses among the observed goods $x_{1},\ldots,x_{p}$ according to the probability vector $A_{{\bm{\theta}}}(x_{j}\mid x_{1},\dots,x_{p})$. This is an appropriate model for the choices of a single customer (or, more generally, for a group of customers with identical preferences). To model the behavior of a large number of consumers with heterogeneous preferences, Boyd and Mellman (1980) and Cardell and Dunbar (1980) suggested to model the population as consisting of a mixture of consumers with different taste vectors. The aggregate probabilities of individual goods being selected is then given by the softmax mixture model (3). This model has been broadly adopted throughout the management science and econometrics literature due to its flexibility and practicality, see (McFadden and Train, 2000; Johnston et al., 2017; Train, 2009; Cameron and Trivedi, 2005) and references therein.
  • Embedded topic modeling. Topic models (Blei et al., 2003) are widely used to represent high-dimensional probability vectors in text analysis, biology, and signal processing; in the standard terminology of the field, each “document” in a corpus is assumed to have been generated as a mixture of “topics”, which are shared across the corpus, each of which is a probability distribution over a large dictionary of “words.” A substantial line of work in statistics and computer science (Arora et al., 2012; Recht et al., 2012; Arora et al., 2013; Bing et al., 2020a, b; Klopp et al., 2023; Wu et al., 2023; Ke and Wang, 2024) has developed conditions under which both the topics and the per-document mixture weights can be consistently estimated from a corpus, at optimal rates. However, since these models do not incorporate any semantic information, the estimated topics can sometimes be difficult to interpret, and the performance of these models can drop dramatically as the dictionary size increases (see, e.g. Dieng et al., 2020). Nguyen et al. (2015) therefore proposed to augment topic model by incorporating semantically meaningful embeddings of each word in the dictionary into a vector space, which can be computed by a number of off-the-shelf approaches in natural language processing and deep learning (Almeida and Xexéo, 2019; Wang et al., 2020). Given embeddings $X_{1},\dots,X_{p}$ of $p$ different words in a dictionary, Nguyen et al. (2015) model each topic as the probability vector $A_{{{\bm{\theta}}}_{k}}(\cdot\mid X_{1},\ldots,X_{p})$, where ${\bm{\theta}}_{k}$ is a latent feature vector associated with the $k$th topic. The topic model assumption then models each document as consisting of words drawn from a multinomial distribution whose probability parameter vector is a softmax mixture (3). Dieng et al. (2020) call this the “embedded topic model,” and show empirically that it improves both interpretability and performance in downstream tasks.
  • Large language models. Almost all modern language models—from some of the earliest neural network approaches (Bengio et al., 2000) to modern architectures based on transformers (Vaswani et al., 2017)—use the softmax function to compute the probability of a given word or token occurring next in a sequence.
    For example, in the paper introducing GPT-1 (Radford et al., 2018), the probability that a word $x_{j}\in\mathbb{R}^{L}$ occurs after a sequence of tokens $u=(u_{1},\dots,u_{m})$, with each $u_{i}\in\mathbb{R}^{L}$, is modeled as $A_{h(u)}(x_{j}\mid X_{1},\dots,X_{p})$, where $X_{1},\dots,X_{p}$ are word embeddings and $h(u)$ is the output of stacked transformer blocks. Although it is not our intention to give a full overview of this setting, for which we refer the reader to Vaswani et al. (2017), we offer a brief commentary that connects it to our framework. Mathematically, a stack of transformer blocks is a composition of functions $h=f_{1}\circ\dots\circ f_{D}$, for some depth $D$, which maps the context $u\in\mathbb{R}^{m\times L}$ to a vector $h(u)$ in $\mathbb{R}^{L}$. The goal of the training process is to estimate these functions so that $h(u)$ is a good prediction of the next token after the sequence $(u_{1},\dots,u_{m})$, in the following sense. The model assumes that the next token comes from a vast library of candidate words, each of which is represented by an embedding vector, $X_{j}\in\mathbb{R}^{L}$, $j\in[p]$. Assuming $h(u)$ is known, the prediction of the next token is given by a probability distribution over the set of $p$ candidate tokens, where the tokens $x_{j}$ are selected with probability given by the softmax vector $A_{h(u)}(x_{j}\mid X_{1},\dots,X_{p})$. In other words, given a fixed context $u$, GPT-1 models the next token via a single multinomial logit, with parameter $h(u)$. Notably, in applications, such as in the GPT family of models (Radford et al., 2018), $p$ is in the tens of thousands, and the embeddings $X_{j}$ are initialized to be i.i.d. Gaussian random variables—this motivates our focus on the large-$p$, $X_{1},\dots,X_{p}$ random setting we consider in this work.
    In statistical language, the “training” process of GPT-1 consists in attempting to estimate the function $h$ from data, which consists of an enormous corpus from books and the internet. However, for a fixed context $u$, it is possible to view this as a parametric estimation problem, with the goal of estimating a single unknown parameter ${\bm{\theta}}:=h(u)$ using a multinomial logit model. Given a large set of $N$ documents in which the same context $(u_{1},\dots,u_{m})$ appears, the “next words” appearing after this common prompt in each of these sentences can be viewed as a sample of size $N$ from a multinomial logit model, whose parameter ${\bm{\theta}}$ can then be estimated by classical techniques.
    The single multinomial logit model described above is common to early large language models, but it can be restrictive: Yang et al. (2017) argued that it fails to capture complex features of natural language. They therefore proposed that language models should model the probability that a word $x_{j}$ occurs after a sequence $u$ as
    $\sum_{k=1}^{K}g_{k}(u)A_{h_{k}(u)}(x_{j}\mid X_{1},\dots,X_{p}),$ (5)
    where $h_{1},\dots,h_{K}$ are $K$ different neural networks, and the coefficients $g_{1},\dots,g_{K}$ are computed by an auxiliary neural network based on the context $u$. In other words, for any given $u$, this is a softmax mixture model (3), with parameters $\alpha_{k}:=g_{k}(u)$ and ${\bm{\theta}}_{k}=h_{k}(u)$. As before, for a fixed context $u$, we can offer of an alternative estimation strategy for $\alpha_{k}$ and ${\bm{\theta}}_{k}$ by estimating parameters of a softmax mixture from the sample of size $N$ obtained as above. This strategy could have wider applicability as, more generally, outputs which take the form of softmax mixtures are characteristic of the “mixture-of-experts” (MoE) approach (Shazeer et al., 2017) in which $A_{h_{k}(u)}(\cdot\mid X_{1},\ldots,X_{p})$ in (5) is the $k$th “expert network” while $g_{k}(u)$ corresponds to the output of its “gating network”. The MoE approach has become an industry default (Fedus et al., 2022; Lepikhin et al., 2020) and is conjectured to be the primary architecture for GPT-4 (OpenAI, 2023).
Throughout, our focus is on estimating the parameters ${\bm{\omega}}^{*}=({\bm{\alpha}}^{*},{\bm{\theta}}_{1}^{*},\dots,{\bm{\theta}}_{K}^{*})$ rather than on the simpler task of estimating the mixture distribution $\pi_{{\bm{\omega}}^{*}}$. Our focus on the parameters is motivated by applications (such as those in discrete choice and topic models) where the parameters reflect important latent structure in the observations (corresponding, for instance, to different customer types or to semantically meaningful topics). Even in the context of large language models, proving rates of estimation of the parameters ${\bm{\omega}}$ for a fixed context provides a theoretical baseline for the error rates of learning the full function $h$ described above, when one fixes the context.
Our bounds on rates of estimation of ${\bm{\omega}}^{*}$ trivially imply corresponding bounds for estimation of $\pi_{{\bm{\omega}}^{*}}$ via the inequality
$\|\pi_{{\bm{\omega}}}-\pi_{{\bm{\omega}}^{\prime}}\|_{1}\leq\|\bm{\alpha}-\bm{\alpha}^{\prime}\|_{1}+\max_{k\in[K]}\max_{j\in[p]}|x_{j}^{\top}({\bm{\theta}}_{k}-{\bm{\theta}}_{k}^{\prime})|\,.$ (6)
Rates of estimation for $\pi_{{\bm{\omega}}^{*}}$ could also be obtained more directly by maximum-likelihood estimation (including via the non-parametric maximum likelihood approach (Kiefer and Wolfowitz, 1956)), which are known to achieve minimax-optimal rates in related settings (see, e.g., Vinayak et al., 2019). Crucially, however, unlike our estimators, it is not known how to implement these estimators in time polynomial in the embedding dimension $L$.
We are interested in a setting where $K$ is a small, fixed constant, but where $L$ is of moderate size, with $L<p$, and $p$ and $N$ are both very large. For example, in the context of language models, $L$ is often taken to be between $500$ and $1000$, whereas $p$ and $N$ can both be in the tens of thousands or larger. A key challenge in this setting is that maximizing the non-convex likelihood corresponding to the model (4) typically requires time scaling exponentially in the dimension of the parameter space (here, $KL+(K-1)$), which is prohibitive in the regimes we consider. The main contribution of this paper is to develop statistically optimal estimators for ${\bm{\omega}}^{\star}$ which are computationally feasible, with runtime polynomial in $L$.
1.1 Our contribution: learning softmax mixtures with many support points
Despite the renewed interest in softmax mixtures, and their ubiquitous usage in AI-type applications, the construction of softmax mixture parameter estimates with provable guarantees has not been studied. In particular, classical estimators, based on the method of moments (MoM) or the Expectation-Maximization (EM) algorithm, have not been shown to provably succeed for softmax mixtures. This motivates our work on:
@@ 1.1 Our contribution: learning softmax mixtures with many support points
    Method of Moments (MoM) parameter estimation in softmax mixtures,

### v2

@@ Abstract
Softmax mixture models (SMMs) are discrete $K$-mixtures introduced to model the probability of choosing an attribute $x_{j}\in\mathbb{R}^{L}$ from $p$ possible candidates, in heterogeneous populations. They have been known, for several decades, as mixed multinomial logits in the econometrics literature, and are gaining traction in the LLM literature, where single softmax models are routinely used in the final layer of a neural network. The theoretical understanding of this mixture model lags behind its growing popularity, and we aim to narrow this gap in this work.
This paper provides a comprehensive analysis of the Expectation-Maximization (EM) algorithm for SMMs, in high dimensions. It complements and extends existing results currently restricted to Gaussian Mixture Models (GMMs). Its population-level theoretical analysis offers key insights into the model that go beyond the typical parameter estimation EM usage. It forms the basis for proving (i) local identifiability, in SSMs with generic features and, further, via a stochastic argument, (ii) full identifiability in SSMs with random features, when $p$ is large enough. To the best of our knowledge, these are the first results in this direction for SSMs with $L>1$.
The population-level EM analysis includes the characterization of the initialization radius for algorithmic convergence. This also guides the construction of possible warm starts of the sample level EM algorithm. Under any warm start initialization, the EM algorithm is shown to recover the mixture atoms of the SSM at the parametric rate, up to logarithmic factors.
We provide two main directions for warm start construction, both based on a new method for estimating the moments of the mixing measure underlying an SSM with random design. First, we construct a method of moments (MoM) preliminary estimator of the mixture parameters, and provide its first theoretical analysis in SSMs. While MoM can enjoy parametric rates of convergence, and thus can serve as a warm-start, the estimator’s quality degrades exponentially in $K$, a fact already demonstrated for GMMs, even when $L=1$. Our recommendation, especially when $K$ is not small, is to follow common practice and run the EM algorithm several times with random initializations. We again make use of the novel estimation method tailored to latent moments in SSMs to further estimate the $K$-dimensional subspace of $\mathbb{R}^{L}$ spanned by the atoms of the mixture. Sampling from this subspace reduces substantially the number of required draws, from $\exp(L)$ to $\exp(K)$, and is also shown to have empirical success.
@@ 1.1 The softmax mixture model
“Softmax mixtures” define a parametric discrete mixture model $\pi\in\Delta^{p}$, the probability simplex in $\mathbb{R}^{p}$, supported on a known set of vectors $x_{1},\dots,x_{p}\in\mathbb{R}^{L}$.
For a given, known and finite $K$, we let ${\bm{\theta}}_{k}\in\mathbb{R}^{L}$, $k\in[K]:=\{1,\ldots,K\}$ be distinct vectors in $\mathbb{R}^{L}$. Each mixture component $A({\bm{\theta}}_{k}):=A(\cdot;{\bm{\theta}}_{k})$ of a softmax mixture is a probability vector in $\Delta^{p}$, supported on $x_{1},\ldots,x_{p}$, parametrized via the softmax function $\mathsf{softmax}:\mathbb{R}^{p}\to\Delta^{p}$,
$A(x_{j};{\bm{\theta}}_{k})=\bigl{[}\mathsf{softmax}(x_{1}^{\top}{\bm{\theta}}_{k},\ldots,x_{p}^{\top}{\bm{\theta}}_{k})\bigr{]}_{j}=\frac{{\exp({x_{j}^{\top}{{\bm{\theta}}_{k}}})}}{\sum_{i=1}^{p}\exp({x_{i}^{\top}{{\bm{\theta}}_{k}}})},$ (1)
for each $j\in[p]$. If we let ${\bm{\alpha}}:=(\alpha_{1},\ldots,\alpha_{K})^{\top}\in\Delta^{K}$ denote the vector of mixing weights, and write ${\bm{\omega}}:=({\bm{\alpha}},{\bm{\theta}}_{1},\ldots,{\bm{\theta}}_{K})$, the softmax mixture model is given by
$\pi(y;{\bm{\omega}}):=\sum_{k=1}^{K}\alpha_{k}A(y;{\bm{\theta}}_{k}),\qquad\mbox{for}\ y\in\{x_{1},\ldots,x_{p}\}.$ (2)
Throughout this paper, our focus is on estimating the parameters ${\bm{\omega}}^{*}=({\bm{\alpha}}^{*},{\bm{\theta}}_{1}^{*},\dots,{\bm{\theta}}_{K}^{*})$ from a sample $Y_{1},\ldots,Y_{N}$ from $\pi^{*}(y):=\pi(y;{\bm{\omega}}^{*})$.
@@
When $K=1$, the softmax mixture model reduces to what is known in the classical statistical literature as the conditional logit model (McFadden, 1974). Its usage and properties, when both $p$ and $L$ are fixed, have been thoroughly studied, see McFadden (1974) and the literature review in (Agresti, 1990, Chapter 9). Much less is known about the case $K>1$, which has received very little attention in the mathematical statistics literature. This paper bridges this gap, and also complements and extends the existing literature on parameter estimation via the expectation-maximization (EM) algorithm beyond the well-studied case of Gaussian Mixture Models with $K$ components ($K$-GMM). We highlight the main contributions of this paper below.
@@
    We develop a hybrid EM algorithm for parameter estimation under softmax mixture models and prove that it converges to the true model parameters at a near-parametric rate after $\mathcal{O}(\log N)$ iterations. Each iteration has computational complexity $\mathcal{O}(pL)$. Our analysis gives conditions on the choice of the algorithm’s initialization, and on the separation between mixture components, under which EM converges. Notably, and improving upon the sharpest known result, albeit developed only for $K$-GMMs, we require that the atom separation depend only logarithmically on the number of components and the smallest mixing weights. As a consequence of the convergence of the population-level EM algorithm, we prove that softmax mixtures are locally identifiable. Section 1.2.1 gives more details and the background for these results, which are formally stated and proved in Sections 2.1, 2.2 and 2.3.